# Grounded GUI Understanding for Vision Based Spatial Intelligent Agent: Exemplified by Virtual Reality Apps

### v3

@@ 6.1.1. RQ1-1: Performance in terms of interactability
As shown in Table 1, Orienter shows consistent performance in interactability detection across different IoU thresholds. Consider the generally best model Orienter with Gemini, at the 0.75 threshold, the mAP peaks at 35.03%, indicating a strong ability to detect IGEs accurately. However, as the IoU threshold increases, indicating a stricter criterion for object detection, there is a noticeable decrease in Precision, dropping to 5.64% at an IoU of 0.95. This trend suggests that while the Orienter is quite effective at a broader detection scope, its precision in highly specific object identification contexts is limited. Orienter demonstrates a balanced performance at lower IoU thresholds, with a peak F1 Score of 46.00% at an IoU of 0.75. This balance is crucial for practical applications, as it indicates a well-rounded capability in both correctly identifying IGEs and minimizing false positives.
For the comparative evaluation, Orienter generally demonstrates superior performance across IoU thresholds when compared with baseline models, especially on the genre split of the dataset, highlighting Orienter’s capability to handle various genres of apps. At an IoU threshold of 0.75, Orienter achieves a Precision of 49.45%, which is approximately 48.0% higher than the best baseline, Faster-RCNN (Ren et al., 2017), at 33.42%. This trend of superiority is more significant at higher IoU thresholds. At the stringent IoU of 0.95, Orienter maintains a Precision of 19.02%, outperforming the closest baseline, CenterNet2 (Zhou et al., 2021), by over 321.7%, a significant margin considering the complexity of VR environments. In terms of Recall and F1 Score, Orienter significantly outperforms other models by considerable margins. At an IoU of 0.75, Orienter’s Recall and F1 Score are 43% and 46.00%, respectively, surpassing CenterNet2 (Zhou et al., 2021)’s 30% and 29.32% by a notable margin of 43.3% and 56.9%, respectively. This indicates that Orienter is not only accurate but also reliable in identifying IGEs, a key requirement in industrial VR apps. Note that the direct prompt method on LMMs performs poorly. This is because although the LMM can identify the objects well, it can not effectively localize them, as discussed in section 2.2. This result again reveals the inability of LMM alone to address the challenge in IGE detection, highlighting Orienter’s enhancement for LMM on this task.
@@ 6.1.2. RQ1-2: Performance in terms of semantics
As for the comparative evaluation in terms of semantics, the results in Table 1 reveal Orienter’s remarkable capability in semantic understanding. At an IoU of 0.75, Orienter achieves a Precision of 23.72% with Gemini on the genre split, which is significantly higher than the best baseline CenterNet2 (Zhou et al., 2021)’s 0.62%, marking an increase of 3725.8%. Similarly, at an IoU of 0.95, Orienter’s F1 Score of 14.31% far surpasses the closest baseline, YOLO v8 (Jocher et al., 2023), which scores 0.40%. This represents an increase of up to 3477.5%, highlighting Orienter’s ability to understand complex semantic structures within VR environments. The direct prompting method continues to perform poorly in terms of semantics due to the aforementioned challenges that LMM alone can not effectively address. Note that the method directly prompting Gemini achieves a recall of 50% on the genre split at 0.75 IoU threshold. This result suggests that LMM can perform well in some special settings.
Our analysis demonstrates that Orienter significantly outperforms existing baseline models in both interactability and semantic understanding. The framework exhibits remarkable precision and recall rates, particularly in high IoU thresholds, which are critical for the nuanced and complex nature of industrial VR apps. This efficacy, especially in semantic understanding, underscores Orienter’s potential in revolutionizing IGE detection in VR environments.
@@ 6.1.3. RQ1-3: Performance in terms of context-sensitive interactability
Methods Interactability Semantics
App Split Genre Split App Split Genre Split
Faster- RCNN mAP18.05 13.84 7.32 1.47 0.07 17.90 12.88 6.96 1.56 0.13 0.21 0.15 0.03 0.01 $\approx$0.0 $\approx$0.0 $\approx$0.0 $\approx$0.0 $\approx$0.0 $\approx$0.0
P33.41 32.29 23.04 11.08 2.15 33.42 29.55 20.10 8.06 1.14 0.91 0.83 0.38 0.16 0.06 0.07 0.06 $\approx$0.0 $\approx$0.0 $\approx$0.0
R32.00 25.00 18.00 8.00 1.00 29.00 22.00 15.00 6.00 1.00 4.00 4.00 3.00 3.00 4.00 2.00 2.00 $\approx$0.0 $\approx$0.0 $\approx$0.0
F1% 32.69 28.18 20.21 9.29 1.36 31.05 25.22 17.18 6.88 1.06 1.48 1.37 0.68 0.30 0.11 0.13 0.12 $\approx$0.0 $\approx$0.0 $\approx$0.0
Center- Net2 mAP17.79 14.98 11.73 5.91 0.93 15.63 12.53 9.07 4.23 0.45 0.20 0.18 0.15 0.05 $\approx$0.0 0.22 0.19 0.05 0.02 $\approx$0.0
P32.53 28.52 26.85 18.26 4.83 28.67 25.81 27.71 16.46 4.51 0.68 0.65 0.57 0.27 0.05 0.62 0.59 0.27 0.19 0.04
R32.00 29.00 23.00 15.00 3.00 30.00 27.00 18.00 12.00 3.00 4.00 4.00 3.00 3.00 2.00 5.00 5.00 3.00 3.00 3.00
Yolo v8 mAP15.30 13.12 9.85 6.16 1.51 11.54 10.02 7.59 3.86 0.46 0.02 0.02 0.01 $\approx$0.0 $\approx$0.0 0.03 0.02 0.02 0.01 0.01
P26.58 25.38 20.63 16.65 7.31 23.92 21.38 24.22 14.55 2.68 0.21 0.20 0.19 0.05 0.01 0.29 0.28 0.27 0.25 0.21
R26.00 23.00 20.00 14.00 3.00 23.00 22.00 15.00 11.00 3.00 4.00 4.00 4.00 1.00 3.00 3.00 3.00 3.00 3.00 3.00
[…]
P0.21 $\approx$0.0 $\approx$0.0 $\approx$0.0 $\approx$0.0 0.43 0.43 $\approx$0.0 $\approx$0.0 $\approx$0.0 0.25 $\approx$0.0 $\approx$0.0 $\approx$0.0 $\approx$0.0 0.23 $\approx$0.0 $\approx$0.0 $\approx$0.0 $\approx$0.0
[…]
F1% $\approx$0.0 $\approx$0.0 $\approx$0.0 $\approx$0.0 $\approx$0.0 $\approx$0.0 $\approx$0.0 $\approx$0.0 $\approx$0.0 $\approx$0.0 0.47 $\approx$0.0 $\approx$0.0 $\approx$0.0 $\approx$0.0 0.46 $\approx$0.0 $\approx$0.0 $\approx$0.0 $\approx$0.0
gpt-4o- 2024-08-06 mAP0.01 $\approx$0.0 $\approx$0.0 $\approx$0.0 $\approx$0.0 0.01 $\approx$0.0 $\approx$0.0 $\approx$0.0 $\approx$0.0 $\approx$0.0 $\approx$0.0 $\approx$0.0 $\approx$0.0 $\approx$0.0 $\approx$0.0 $\approx$0.0 $\approx$0.0 $\approx$0.0 $\approx$0.0
P0.78 $\approx$0.0 $\approx$0.0 $\approx$0.0 $\approx$0.0 0.75 $\approx$0.0 $\approx$0.0 $\approx$0.0 $\approx$0.0 $\approx$0.0 $\approx$0.0 $\approx$0.0 $\approx$0.0 $\approx$0.0 $\approx$0.0 $\approx$0.0 $\approx$0.0 $\approx$0.0 $\approx$0.0
[…]
F1% $\approx$0.0 $\approx$0.0 $\approx$0.0 $\approx$0.0 $\approx$0.0 $\approx$0.0 $\approx$0.0 $\approx$0.0 $\approx$0.0 $\approx$0.0 $\approx$0.0 $\approx$0.0 $\approx$0.0 $\approx$0.0 $\approx$0.0 $\approx$0.0 $\approx$0.0 $\approx$0.0 $\approx$0.0 $\approx$0.0
Orienter (gpt-4-vision- preview) mAP25.53 24.32 23.03 19.64 4.61 27.07 25.22 23.87 19.80 4.54 14.93 14.66 13.96 12.17 3.67 15.96 15.57 14.62 12.87 3.75
P34.79 33.52 32.32 29.31 14.92 38.42 36.70 34.97 31.40 16.79 16.06 16.05 15.17 15.27 6.53 17.11 16.85 15.78 15.93 6.14
R40.00 39.00 38.00 34.00 12.00 40.00 38.00 37.00 32.00 11.00 50.00 50.00 50.00 33.00 20.00 50.00 50.00 50.00 33.00 20.00
Orienter (gpt-4o- 2024-08-06) mAP26.44 25.26 23.98 20.35 4.48 26.11 24.50 23.51 19.42 4.38 18.00 17.65 16.91 14.28 5.75 19.45 18.97 18.07 15.40 5.68
P34.63 34.89 35.43 31.64 18.57 37.91 36.68 36.71 33.24 12.91 19.32 18.97 18.28 15.37 6.87 20.35 19.81 19.01 16.19 7.36
R42.00 39.00 36.00 34.00 11.00 41.00 39.00 37.00 33.00 14.00 50.00 50.00 50.00 50.00 42.00 50.00 50.00 50.00 50.00 25.00
Orienter (Claude 3.5 Sonnet) mAP23.99 22.90 21.54 18.46 3.90 25.50 23.96 22.21 18.49 3.84 16.14 15.79 15.17 12.73 5.01 19.01 18.51 17.78 15.00 6.67
P31.73 37.52 36.79 33.56 16.30 37.97 40.15 39.22 35.86 14.51 16.63 16.28 15.83 13.35 7.64 19.68 19.11 18.53 15.75 9.87
R40.00 32.00 31.00 29.00 11.00 37.00 33.00 32.00 29.00 13.00 50.00 50.00 50.00 50.00 20.00 50.00 50.00 50.00 50.00 20.00
Orienter (Gemini 1.5 Pro) mAP32.71 31.27 30.51 26.33 5.88 35.03 32.59 30.82 25.08 5.64 20.69 20.35 20.19 17.77 7.16 22.12 21.44 21.17 18.45 7.41
P46.42 45.30 44.94 41.64 21.46 49.45 47.52 46.42 41.94 19.02 22.64 22.29 22.16 19.39 9.66 23.72 22.86 22.64 19.58 8.35
R42.00 41.00 40.00 37.00 14.00 43.00 42.00 41.00 37.00 15.00 50.00 50.00 50.00 50.00 25.00 50.00 50.00 50.00 50.00 50.00
    The highest metrics are colored with gray. Metrics below 0.01% are regarded as $\approx$ 0.0. All metrics of UIED and Xianyu are lower than 0.01% and omitted.
@@
Faster-RCNN CenterNet2 Yolo v8 gpt-4o-2024-08-06 Gemini 1.5 Pro
@@
IoU PRF1% PRF1% PRF1% PRF1% PRF1% PRF1% PRF1% PRF1% PRF1%
0.75 2.44 0.14 0.27 4.27 0.27 0.51 $\approx$0.0 $\approx$0.0 $\approx$0.0 2.44 0.07 0.13 $\approx$0.0 $\approx$0.0 $\approx$0.0 44.72 6.66 10.77 67.70 24.00 32.51 62.74 18.54 26.07 67.91 22.96 31.40
0.80 2.44 0.09 0.18 4.88 0.27 0.52 $\approx$0.0 $\approx$0.0 $\approx$0.0 2.44 0.07 0.13 $\approx$0.0 $\approx$0.0 $\approx$0.0 44.57 6.32 10.41 67.99 23.36 31.82 62.93 17.94 25.36 67.91 22.20 30.63
0.85 $\approx$0.0 $\approx$0.0 $\approx$0.0 4.88 0.23 0.43 $\approx$0.0 $\approx$0.0 $\approx$0.0 $\approx$0.0 $\approx$0.0 $\approx$0.0 $\approx$0.0 $\approx$0.0 $\approx$0.0 45.02 6.27 10.40 67.82 22.21 30.62 64.17 16.84 24.27 68.42 21.40 29.75
0.90 $\approx$0.0 $\approx$0.0 $\approx$0.0 4.88 0.18 0.34 $\approx$0.0 $\approx$0.0 $\approx$0.0 $\approx$0.0 $\approx$0.0 $\approx$0.0 $\approx$0.0 $\approx$0.0 $\approx$0.0 45.24 5.37 9.26 68.93 19.35 27.75 63.96 14.84 21.84 70.01 18.23 26.78
0.95 $\approx$0.0 $\approx$0.0 $\approx$0.0 $\approx$0.0 $\approx$0.0 $\approx$0.0 $\approx$0.0 $\approx$0.0 $\approx$0.0 $\approx$0.0 $\approx$0.0 $\approx$0.0 $\approx$0.0 $\approx$0.0 $\approx$0.0 26.83 1.62 3.03 47.56 8.99 13.59 47.07 7.47 11.37 53.35 7.01 11.76
@@
6.2. RQ2: Ablation Study
Table 3 shows the performance of Orienter and its three variations. For Orienter w/o Context Comprehension, we remove the whole Module I Semantic Context Comprehension. For Orienter w/o Reflection-Directed Loop, we remove the whole Module II instead of IGE candidate detection module (to at least localize elements). For Orienter w/o interactability Classification, we remove the whole Module III Context-Sensitive Interactability Classification.
@@ 6.2. RQ2: Ablation Study
6.3. RQ3: Usefulness
@@
Figure 14. Results of RQ3: Usefulness
As shown in Figure 14, the testing guided by Orienter outperforms the random testing with a significant gap. We select the best model, Orienter with Gemini for comparison. The number of effective interactions in guided testing increases notably faster than in random generation. At the end of the testing, the guided one’s average total number of effective interactions reaches 6,571.8, surpassing the random testing’s 2,911.6 by a remarkable 125.7%. Regarding the rate of effective interaction, guided testing peaks at 0.57, with an average of 0.46 over the whole duration of testing, considerably better than that of the random testing remaining at around 0.11. At 10 minutes, the IGE coverage of the guided testing already reaches 0.65, outperforming the random testing’s 0.32 by 103.1%. The IGE coverage of the guided testing eventually reaches 0.83, which is 27.7% higher than the random testing’s 0.65.
These results demonstrate the capability of Orienter to boost the effectiveness and efficiency of VR app testing. Apart from automated VR app testing, other downstream software engineering problems like UI to code, video-based or screenshot-based testing, etc., can also benefit from the ability of Orienter to effective and efficient IGE detection.
@@ 7. Threats to Validity
Internal Validity. An internal threat to validity is related to the annotators’ bias and subjective interpretation. Given that the dataset was annotated by 13 annotators, there’s potential bias based on their personal experience and interpretation of interactability in VR apps. This could impact the consistency and accuracy of the annotations. The process of labeling semantics of IGEs is subjective and dependent on the annotator’s understanding and perception, which may not always align with the intended use or perception in different user demographics. To mitigate it, we construct a detailed and objective data collection and annotation process. We spend at least two hours for each annotator to make sure they understand all steps. Another threat is model bias in pretrained models, given that Orienter relies on pretrained LMMs, there is a risk that biases inherent in these models could affect the accuracy of IGE detection. If these models were trained on data that was not representative of diverse contexts, their predictions could be skewed. To mitigate this threat, we choose the most powerful and representative models, which are trained on large corpus and which are claimed that the team has reduced model bias to some level.
External Validity. One threat to external validity is the generalizability across VR apps. The diversity in VR apps poses a challenge to generalizability. Orienter’s effectiveness demonstrated in specific scenarios may not uniformly apply across different VR environments, particularly those with unique interaction paradigms or novel GUI element types. To mitigate this threat, we sample popular and complicated VR apps, crossing a large portion of app categories.

### v4

@@ 6.1.1. RQ1-1: Performance in terms of interactability
As shown in Table 2, Orienter shows consistent performance in interactability detection across different IoU thresholds. Consider the generally best model Orienter with Gemini, at the 0.75 threshold, the mAP peaks at 35.03%, indicating a strong ability to detect IGEs accurately. However, as the IoU threshold increases, indicating a stricter criterion for object detection, there is a noticeable decrease in Precision, dropping to 5.64% at the IoU of 0.95. This trend suggests that while the Orienter is quite effective at a broader detection scope, its precision in highly specific object identification contexts is limited. Orienter demonstrates a balanced performance at lower IoU thresholds, with a peak F1 Score of 46.00% at an IoU of 0.75. This balance is crucial for practical applications, as it indicates a well-rounded capability in both correctly identifying IGEs and minimizing false positives.
For the comparative evaluation, Orienter generally demonstrates superior performance across IoU thresholds when compared with baseline models, especially on the genre split of the dataset, highlighting Orienter’s capability to handle various genres of apps. At an IoU threshold of 0.75, Orienter achieves a Precision of 49.45%, which is approximately 48.0% higher than the best baseline, Faster-RCNN (Ren et al., 2017), at 33.42%. This trend of superiority is more significant at higher IoU thresholds. At the stringent IoU of 0.95, Orienter maintains a Precision of 19.02%, outperforming the closest baseline, CenterNet2 (Zhou et al., 2021), by over 321.7%, a significant margin considering the complexity of XR environments. In terms of Recall and F1 Score, Orienter significantly outperforms other models by considerable margins. At an IoU of 0.75, Orienter’s Recall and F1 Score are 43% and 46.00%, respectively, surpassing CenterNet2 (Zhou et al., 2021)’s 30% and 29.32% by a notable margin of 43.3% and 56.9%, respectively. This indicates that Orienter is not only accurate but also reliable in identifying IGEs, a key requirement in industrial XR apps. Note that the direct prompt method on LMMs performs poorly. This is because although the LMM can identify the objects well, it can not effectively localize them, as discussed in section 2.1. This result again reveals the inability of LMM alone to address the challenge in IGE detection, highlighting Orienter’s enhancement for LMM on this task.
@@ 6.1.2. RQ1-2: Performance in terms of semantics
As for the comparative evaluation in terms of semantics, the results in Table 2 reveal Orienter’s remarkable capability in semantic understanding. At an IoU of 0.75, Orienter achieves a Precision of 23.72% with Gemini on the genre split, which is significantly higher than the best baseline CenterNet2 (Zhou et al., 2021)’s 0.62%, marking an increase of 3725.8%. Similarly, at an IoU of 0.95, Orienter’s F1 Score of 14.31% far surpasses the closest baseline, YOLO v8 (Jocher et al., 2023), which scores 0.40%. This represents an increase of up to 3477.5%, highlighting Orienter’s ability to understand complex semantic structures within XR environments. The direct prompting method continues to perform poorly in terms of semantics due to the aforementioned challenges that LMM alone can not effectively address. Note that the method directly prompting Gemini achieves a recall of 50% on the genre split at 0.75 IoU threshold. This result suggests that LMM can perform well in some special settings.
Our analysis demonstrates that Orienter significantly outperforms existing baseline models in both interactability and semantic understanding. The framework exhibits remarkable precision and recall rates, particularly in high IoU thresholds, which are critical for the nuanced and complex nature of industrial XR apps. This efficacy, especially in semantic understanding, underscores Orienter’s potential in revolutionizing IGE detection in XR environments.
@@ 6.1.3. RQ1-3: Performance in terms of context-sensitive interactability
Faster-RCNN CenterNet2 Yolo v8 OmniParser gpt-4o-2024-08-06
IoU P% R% F1% P% R% F1% P% R% F1% P% R% F1% P% R% F1%
0.75 2.44 0.14 0.27 4.27 0.27 0.51 ¡0.01 ¡0.01 ¡0.01 ¡0.01 ¡0.01 ¡0.01 2.44 0.07 0.13
0.80 2.44 0.09 0.18 4.88 0.27 0.52 ¡0.01 ¡0.01 ¡0.01 ¡0.01 ¡0.01 ¡0.01 2.44 0.07 0.13
0.85 ¡0.01 ¡0.01 ¡0.01 4.88 0.23 0.43 ¡0.01 ¡0.01 ¡0.01 ¡0.01 ¡0.01 ¡0.01 ¡0.01 ¡0.01 ¡0.01
0.90 ¡0.01 ¡0.01 ¡0.01 4.88 0.18 0.34 ¡0.01 ¡0.01 ¡0.01 ¡0.01 ¡0.01 ¡0.01 ¡0.01 ¡0.01 ¡0.01
0.95 ¡0.01 ¡0.01 ¡0.01 ¡0.01 ¡0.01 ¡0.01 ¡0.01 ¡0.01 ¡0.01 ¡0.01 ¡0.01 ¡0.01 ¡0.01 ¡0.01 ¡0.01
Gemini 1.5 Pro
@@
IoU P% R% F1% P% R% F1% P% R% F1% P% R% F1% P% R% F1%
0.75 ¡0.01 ¡0.01 ¡0.01 44.72 6.66 10.77 67.70 24.00 32.51 62.74 18.54 26.07 67.91 22.96 31.40
0.80 ¡0.01 ¡0.01 ¡0.01 44.57 6.32 10.41 67.99 23.36 31.82 62.93 17.94 25.36 67.91 22.20 30.63
0.85 ¡0.01 ¡0.01 ¡0.01 45.02 6.27 10.40 67.82 22.21 30.62 64.17 16.84 24.27 68.42 21.40 29.75
0.90 ¡0.01 ¡0.01 ¡0.01 45.24 5.37 9.26 68.93 19.35 27.75 63.96 14.84 21.84 70.01 18.23 26.78
0.95 ¡0.01 ¡0.01 ¡0.01 26.83 1.62 3.03 47.56 8.99 13.59 47.07 7.47 11.37 53.35 7.01 11.76
6.2. RQ2: Ablation Study
@@
Table 4 shows the performance of Orienter and its three variations. We remove the corresponding components of each variation to demonstrate their neceesity.
@@
6.3. RQ3: Effectiveness for boosting automated testing
Table 5. Average testing result by apps with different categories. Categories with # of app ¡ 10 are omitted
App Category # App # GUI Scene Random IGE Cov. Guided IGE Cov. Improved
All 77 936 0.65 0.83 +26.91%
Singleplayer 43 640 0.45 0.60 +32.42%
Casual 46 600 0.42 0.52 +24.75%
Simulation 30 521 0.37 0.49 +31.05%
3D 24 471 0.35 0.47 +33.43%
Immersive Sim 13 374 0.28 0.38 +35.05%
First-Person 28 402 0.29 0.36 +25.61%
Puzzle 14 339 0.25 0.34 +36.60%
Free to Play 36 393 0.27 0.33 +20.64%
Early Access 11 319 0.24 0.32 +32.52%
Indie 38 407 0.27 0.32 +19.88%
Colorful 13 318 0.24 0.32 +35.39%
6DOF 14 295 0.21 0.27 +29.22%
Education 12 292 0.21 0.27 +31.40%
Atmospheric 24 267 0.18 0.21 +15.05%
Adventure 23 224 0.15 0.19 +26.29%
Action 24 220 0.15 0.17 +13.06%
Physics 14 178 0.12 0.15 +21.87%
Exploration 14 181 0.12 0.15 +21.17%
Arcade 17 163 0.11 0.13 +25.59%
Funny 10 93 0.05 0.07 +34.46%
Strategy 11 87 0.06 0.07 +13.52%
Sci-fi 11 65 0.04 0.04 +8.84%
Multiplayer 11 56 0.03 0.03 -11.37%
Mean 20 300 0.21 0.27 +24.19%
@@ 6.3. RQ3: Effectiveness for boosting automated testing
Figure 15. Performance of testing over time across strategies
As shown in Figure 15, the testing guided by Orienter outperforms the random testing with a significant gap. We select the best model, Orienter with Gemini for comparison. The number of effective interactions in guided testing increases notably faster than in random generation. At the end of the testing, the guided one’s average number of effective interactions reaches 13,038.03 on 936 GUI Scenes in total, surpassing the random testing’s 5,911.53 by a remarkable 120.6%. At 10 minutes, the IGE coverage of the guided testing already reaches 0.64, surpassing the random testing’s 0.30 by 113.3%. The guided testing eventually reaches a coverage of 0.83, 27.7% higher than the random testing’s 0.65.
Figure 16. Coverage rate on apps in different categories. Categories with # of app ¡ 10 are omitted
We further investigate the performance of guided and non-guided testing agents on apps with different categories. We select 23 community-generated categories with no less than 10 apps (except for the category ”VR”) for category-wise evaluation in terms of IGE coverage, covering 76 of the 77 AUTs. Table 5 and Figure 16 intuitively reveal the benefits of the guidance from Orienter. The testing agent with the guidance of Orienter surpasses the one without guidance in most app categorys, achieving improvements by up to 36.60% (0.34 vs. 0.25) in the category ”Puzzle”. The guided agent achieves the highest IGE coverage of 0.60 in the category ”Singleplayer”, surpassing the random agent by 32.42%. In contraction, the guided testing agent shows a worse yet still very close IGE coverage compared with the random testing agent in the ”Multiplayer” category, both achieving a coverage at around 0.03. Such reduction suggests that the bias introduced by the external guidance could also lead to performance degradation for testing agents, highlighting the necessity of ”thinking outside the box” to explore the GUI scene aside from the guidance to some extent. The large gap between the two opposite categories also reflects the huge difference between different categories of apps. In addition, the guided testing agent generally performs more consistently than the random testing agent between testing rounds.
These results demonstrate the capability of Orienter to boost the effectiveness and efficiency of XR app testing. Apart from automated XR app testing, other downstream software engineering problems like UI to code, video-based or screenshot-based testing, etc., can also benefit from the ability of Orienter to effective and efficient IGE detection.
@@ 7. Threats to Validity
Internal Validity. An internal threat to validity is related to the annotators’ bias and subjective interpretation. Given that the dataset was annotated by 13 annotators, there’s potential bias based on their personal experience and interpretation of interactability in XR apps. This could impact the consistency and accuracy of the annotations. The process of labeling semantics of IGEs is subjective and dependent on the annotator’s understanding and perception, which may not always align with the intended use or perception in different user demographics. To mitigate it, we construct a detailed and objective data collection and annotation process. We spend at least two hours for each annotator to make sure they understand all steps. Another threat is model bias in pretrained models, given that Orienter relies on pretrained LMMs, there is a risk that biases inherent in these models could affect the accuracy of IGE detection. If these models were trained on data that was not representative of diverse contexts, their predictions could be skewed. To mitigate this threat, we choose the most powerful and representative models, which are trained on large corpus and which are claimed that the team has reduced model bias to some level.
External Validity. One threat to external validity is the generalizability across XR apps. The diversity in XR apps poses a challenge to generalizability. Orienter’s effectiveness demonstrated in specific scenarios may not uniformly apply across different XR environments, particularly those with unique interaction paradigms or novel GUI element types. To mitigate this threat, we sample popular and complicated XR apps, crossing a large portion of app categories.